OpenStudy (anonymous):

Graph each parabola and list the vertex for each. y = x^2 - 4

OpenStudy (anonymous):

http://celestia.00freehost.com

OpenStudy (anonymous):

What...

OpenStudy (xapproachesinfinity):

y=x^2-4 is just y=x^2 moved downwards 4 units

OpenStudy (anonymous):

Can you show me how to graph it? I got that part just don't know how to graph. @xapproachesinfinity

OpenStudy (xapproachesinfinity):

do you know how to graph y=x^2

OpenStudy (xapproachesinfinity):

so?

OpenStudy (anonymous):

I figured it out. Thanks man

OpenStudy (xapproachesinfinity):

no prob